OpenStudy (anonymous):

Please help me solve for n... 16n^2-9=0

OpenStudy (anonymous):

\[(4n-3)(4n+3)\\ 4n-3=0\\\\ 4n+3=0\\\\ n= -\frac{3}{4}\\\\ n= \frac{3}{4}\\\\ \]